OpenStudy (anonymous):

a 8000kg tow truck pulls a 3000kg car horizontally into a garage. if the motor of the tow truck exerts a forward thrust on the truck of 800N, what is the tension in the cable connecting the truck to the car? (neglect any forces of friction)

OpenStudy (anonymous):

 m1 : m1a =T (1) m2 : m2a= F - T (2) 1+2 a=0.073 (1) we get T = 218.18N